The Evanec Company's next expected dividend, D1, is $2.89; its growth rate is 4%; and its common stock now sells for $34.00. New stock (external equity) can be sold to net $27.20 per share.

a. What is Evanec's cost of retained earnings, rs? Do not round intermediate calculations. Round your answer to two decimal places.

b. What is Evanec's percentage flotation cost, F? Round your answer to two decimal places.

c. What is Evanec's cost of new common stock, re? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

a.cost of retained earning=(D1/Current price)+Growth rate

=(2.89/34)+0.04

=12.5%

b.percentage flotation cost=(Current price-Net price)/Current price

=(34-27.2)/34

=20%

c. cost of new common stock=(D1/Net price)+Growth rate

=(2.89/27.2)+0.04

=14.63%(Approx).